# Enhancing Pre-Trained Generative Language Models with Question Attended Span Extraction on Machine Reading Comprehension

### v1

@@ Abstract
Machine Reading Comprehension (MRC) poses a significant challenge in the field of Natural Language Processing (NLP). While mainstream MRC methods predominantly leverage extractive strategies using encoder-only models such as BERT, generative approaches face the issue of out-of-control generation – a critical problem where answers generated are often incorrect, irrelevant, or unfaithful to the source text. To address these limitations in generative models for MRC, we introduce the Question-Attended Span Extraction (QASE) module. Integrated during the fine-tuning phase of pre-trained generative language models (PLMs), QASE significantly enhances their performance, allowing them to surpass the extractive capabilities of advanced Large Language Models (LLMs) such as GPT-4. Notably, these gains in performance do not come with an increase in computational demands. The efficacy of the QASE module has been rigorously tested across various datasets, consistently achieving or even surpassing state-of-the-art (SOTA) results. Our code is available at this anonymous repo link.
@@ 1 Introduction
Machine Reading Comprehension (MRC), also referred to as text-grounded question answering (QA) Wang et al. (2022), involves presenting a model with a text passage and a question, requiring it to formulate an answer based solely on the given text. This can be achieved either by identifying a specific span within the text or by generating a concise answer. MRC poses a significant challenge within the domain of Natural Language Processing (NLP). Predominant strategies for addressing MRC employ extractive methods, which typically extract pertinent text snippets from a broader context in response to a query Wang et al. (2018); Yan et al. (2019); Chen et al. (2020). However, the most precise answers in practical settings often span multiple text passages or necessitate inferential reasoning that extends beyond the surface-level content Li et al. (2021). Therefore, there is a compelling necessity to integrate generative models alongside extractive approaches to enhance the robustness and comprehensiveness of solutions in this field.
@@
Yet, generative models often fall short in MRC tasks due to a phenomenon known as out-of-control generation Li et al. (2021), which encompasses two primary issues, as illustrated in Figure 1: (a) ill-formed generations that include incomplete or redundant phrases, and (b) factual inconsistencies that diverge from the intended information. To tackle these issues, this paper introduces the lightweight Question-Attended Span Extraction (QASE) module. This module is integrated during the fine-tuning of various open-source generative pre-trained language models (PLMs) across multiple MRC datasets to enhance the reliability and accuracy of the generated answers.
@@
    We develop the QASE module to enhance the quality and factual accuracy of answers generated by fine-tuned generative PLMs, achieving performance on par with state-of-the-art (SOTA) extractive methods and surpassing that of advanced Large Language Models (LLMs) such as GPT-4.
@@
    QASE enhances model performance without imposing significant additional computational demands, offering a cost-effective solution for researchers operating under resource constraints.
@@ 2 Related Work
Research in MRC
In recent research on MRC, there is a predominant focus on extractive question answering using encoder-only PLMs such as BERT and XLM-Roberta. These studies typically involve predicting the start and end positions of answers directly from the context provided Ohsugi et al. (2019); Lan et al. (2019); Bachina et al. (2021); Chen et al. (2022). Additionally, to accommodate scenarios where answers comprise multiple spans from the text – termed as the multi-span setting – researchers have proposed different strategies. Segal et al. (2020) treated this as a sequence tagging task where each token in the sequence is tagged as being part of an answer span or not. Others Hu et al. (2019); Lee et al. (2023); Zhang et al. (2023) have experimented with hybrid approaches that combine different modeling techniques or tasks to enhance performance on complex MRC problems. Beyond these extractive methods, there is an emerging interest in applying generative language models(GLMs) for MRC Yang et al. (2020); Li et al. (2021); Jiang et al. (2022); Su et al. (2022). These models do not just predict the location of an answer within a given text but generate an answer reformulating information found across the context.
@@
In the domain of controllable text generation, significant progress has been achieved. Gururangan et al. (2020) employ fine-tuning of language models on domain-adaptive text to customize the attributes of generated content. Other promising methodologies include reinforcement learning Li et al. (2024), contrastive learning Zheng et al. (2023), and the use of control codes for fine-tuning PLMs Keskar et al. (2019). Some approaches involve modifying the probability distribution of PLMs. For instance, Liu et al. (2021) propose a technique using two smaller "expert" models to refine the PLM’s output, while Yang and Klein (2021) introduce a method that conditions the generation process using a "future discriminator" based on predicted outcomes. Moreover, Huang et al. (2023) explore efficient multi-aspect text generation with trainable gates for enhanced control. QASE represents a novel adaptation of controlled text generation tailored to the specific challenges of MRC, with a focus on the precision and relevance of generated answers. Unlike methods that modify the overall generative process through complex architectural alterations or additional learning mechanisms, QASE directly utilizes the question context to guide the extraction and generation phases.
@@ 4.1 Datasets and Metrics
We utilize three MRC benchmark datasets:
@@ 4.2 Experimental Setup
4.3 Experiment Results
@@
To assess the QASE module, we compare the performance of various PLMs fine-tuned with and without QASE, as detailed in Table 2. Overall, models fine-tuned with QASE consistently outperform those without it. Specifically, for the SQuAD dataset, models with QASE show an exact match (EM) percentage increase of up to 33.8% and an F1 score improvement of up to 8.4% compared to vanilla fine-tuned models. For MultiSpanQA, improvements include up to 1.6% in EM F1 and up to 3.3% in overlap F1. Likewise, on the Quoref dataset, enhancements of up to 19.2% in EM percentage and up to 16.0% in F1 score are observed. These results confirm that QASE enables generative-based PLMs to produce more accurate, contextually coherent, and higher-quality answers in MRC tasks compared to vanilla fine-tuning approaches.
@@
4.4 Model Comparisons
Our top model, Flan-T5-LargeQASE, is further benchmarked against leading models on each dataset’s official leaderboard, alongside zero-shot GPT-3.5-Turbo and GPT-4. GPT-3.5-Turbo stands as one of OpenAI’s most efficient models in terms of capability and cost, while GPT-4 shows superior reasoning abilities Liu et al. (2023c). Studies indicate their superiority over traditional fine-tuning methods in most logical reasoning benchmarks Liu et al. (2023a). The prompts used to query the GPT variants are detailed in Appendix A.4.
@@ 4.4 Model Comparisons
On SQuAD, as showed in Table 3, Flan-T5-LargeQASE surpasses human performance, equaling the NLNet model from Microsoft Research Asia and the original pre-trained BERT-Large from Google Devlin et al. (2019). Additionally, it surpasses GPT-4 by 113.8% on the exact match score and 32.6% on F1.
@@
On MultiSpanQA, Table 4 shows that Flan-T5-LargeQASE outperforms LIQUID Lee et al. (2023), which currently ranks #1 on the leaderboard, with respect to the overlap F1 score. Moreover, it surpasses GPT-4 by 4.5% on the exact match F1 and 1.5% on the overlap F1.
@@
On Quoref, Table 5 shows that Flan-T5-LargeQASE is comparable to CorefRoberta-Large Ye et al. (2020), which ranks #9 on the leaderboard, with a 0.5% higher exact match. Furthermore, it outperforms GPT-4 by 11.9% on the exact match and 4.8% on F1.
All top-performing models on these datasets’ leaderboards, equaling or exceeding Flan-T5-LargeQASE, are encoder-only extractive models. Therefore, these results demonstrate that QASE-enhanced generative PLMs can be fine-tuned to match or exceed the capabilities of SOTA extractive models and outperform leading LLMs in MRC.
@@ 4.5 Ablation Studies
4.6 Computational Cost
4.7 Factual Consistency
Flan-T5-Large $\bm{Q^{2}}$ F1 $\bm{Q^{2}}$ NLI
SQuAD no QASE 42.927 44.983
QASE 43.624 45.419
MultiSpanQA no QASE 32.889 31.433
QASE 34.732 36.452
4.7.1 Qualitative Case Studies
In addition to the $Q^{2}$ statistical analysis, we also perform qualitative case studies to further demonstrate the effectiveness of QASE in generating factual consistent answers.
@@ 5 Discussions
In this section, we briefly address the weak performance of Flan-T5 zero-shot and Llama 2 on MRC tasks, despite their strong language understanding abilities. We note that a comprehensive analysis is beyond our study’s scope. Our goal is to gain insights into further improving these PLMs’ effectiveness in MRC.
@@ 5.2 Llama 2 Performance
We also observe that models based on Llama 2 and Alpaca consistently underperform compared to those based on Flan-T5, across zero-shot and fine-tuned scenarios, with or without QASE. This discrepancy may arise from the significant difference in the number of trainable parameters, as illustrated in Table 1, during fine-tuning. Additionally, factors such as differences in pre-training datasets and varied adaptation to tasks due to structural disparities can also contribute to this performance gap. While acknowledging these factors, conducting a comprehensive comparison of different generative model architectures in MRC tasks exceeds the scope of our study. For further discussion, please refer to Appendix B.2.
5.3 Prompting Strategies
While we acknowledge the prevalent use of prompting strategies like in-context learning, our main focus is not on evaluating these strategies but rather on improving the performance of generative PLMs on MRC tasks through fine-tuning, with and without the integration of QASE. However, we do experiment with various prompt templates, as discussed in Section 4.5 and Appendix B.1.
@@ 6 Conclusion and Future Work
In this study, we address out-of-control generation issue of generative PLMs in MRC using QASE, a lightweight question-attended span extraction module, during the fine-tuning of PLMs. Our experiments show that QASE-enhanced PLMs generate better-quality responses with improved formality and factual consistency, matching SOTA extractive models and outperforming GPT-4 by a significant margin on all three MRC datasets. Importantly, QASE improves performance without a significant increase in computational costs, benefiting researchers with limited resources.
In the future, we aim to evaluate our model on generative MRC datasets Nguyen et al. (2016) to gauge its effectiveness in handling more intricate scenarios. Additionally, a significant emphasis will be placed on assessing the model’s overall capability in answer generation, with a specific focus on human perception. This involves incorporating human annotators alongside automatic metrics. Looking further ahead, we aspire to extend our research to explore strategies for mitigating input- and context-conflicting hallucinations in LLMs.
@@ A.4 Instruction Templates and Model Prompts
Table 13 provides the instruction and prompt templates used for fine-tuning the PLMs and for zero-shot querying of PLMs and GPT variants across both single- and multi-span answer datasets.
@@
Zero-shot prompting PLMs and GPT variants on single-span answer dataset, SQuAD Instruction: Using the provided context, answer the question with exact phrases and avoid explanations.
@@
Zero-shot prompting PLMs and GPT variants on multi-span answer datasets, MultiSpanQA and Quoref Instruction: Using the provided context, answer the question with exact phrases and avoid explanations. Format the response as follows: ["answer1", "answer2", …].
@@
Table 13: Templates for fine-tuning instructions and zero-shot query prompts

### v2

@@ Abstract
Machine Reading Comprehension (MRC) poses a significant challenge in the field of Natural Language Processing (NLP). While mainstream MRC methods predominantly leverage extractive strategies using encoder-only models such as BERT, generative approaches face the issue of out-of-control generation – a critical problem where answers generated are often incorrect, irrelevant, or unfaithful to the source text. To address these limitations in generative models for extractive MRC, we introduce the Question-Attended Span Extraction (QASE) module. Integrated during the fine-tuning phase of pre-trained generative language models (PLMs), QASE significantly enhances their performance, allowing them to surpass the extractive capabilities of advanced Large Language Models (LLMs) such as GPT-4 in few-shot settings. Notably, these gains in performance do not come with an increase in computational demands. The efficacy of the QASE module has been rigorously tested across various datasets, consistently achieving or even surpassing state-of-the-art (SOTA) results, thereby bridging the gap between generative and extractive models in extractive MRC tasks. Our code is available at this anonymous repo link.
@@ 1 Introduction
Extractive Machine Reading Comprehension (MRC), also referred to as text-grounded question answering (QA) Wang et al. (2022), involves presenting a model with a text passage and a question, requiring it to formulate an answer based solely on the given text. This can be achieved either by identifying a specific span within the text or by generating a concise answer. Extractive MRC poses a significant challenge within the domain of Natural Language Processing (NLP). Predominant strategies for addressing extractive MRC employ extractive methods, which typically extract pertinent text snippets from a broader context in response to a query Wang et al. (2018); Yan et al. (2019); Chen et al. (2020). However, the most precise answers in practical settings often span multiple text passages or necessitate inferential reasoning that extends beyond the surface-level content Li et al. (2021). Therefore, there is a compelling necessity to integrate generative models alongside extractive approaches to enhance the robustness, versatility, and comprehensiveness of solutions in this field.
@@
Yet, generative models often fall short in extractive MRC tasks due to a phenomenon known as out-of-control generation Li et al. (2021), which encompasses two primary issues, as illustrated in Figure 1: (a) ill-formed generations that include incomplete or redundant phrases, and (b) factual inconsistencies that diverge from the intended information. Our research aim to bridge the performance gap between generative and extractive models in extractive MRC tasks by tackling the out-of-control generation issue. We introduce the lightweight Question-Attended Span Extraction (QASE) module. This module is integrated during the fine-tuning of various open-source generative pre-trained language models (PLMs) across multiple MRC datasets to enhance the reliability and accuracy of the generated answers.
@@
    We develop the QASE module to enhance the quality and factual accuracy of answers generated by fine-tuned generative PLMs, achieving performance on par with state-of-the-art (SOTA) extractive methods and surpassing that of advanced Large Language Models (LLMs) such as GPT-4 in few-shot settings.
@@
    QASE enhances model performance without imposing significant additional computational demands, offering a cost-effective solution.
@@ 2 Related Work
Extractive MRC
Recent MRC research predominantly focuses on extractive question answering using encoder-only PLMs like BERT and XLM-Roberta, predicting the start and end positions of answers directly from the context Ohsugi et al. (2019); Lan et al. (2019); Bachina et al. (2021); Chen et al. (2022). For multi-span answers, Segal et al. (2020) treat this as a sequence tagging task, while others Hu et al. (2019); Lee et al. (2023); Zhang et al. (2023) use hybrid approaches to enhance performance on complex MRC problems. Beyond extractive methods, there is growing interest in applying generative language models for extractive MRC Yang et al. (2020); Li et al. (2021); Jiang et al. (2022); Su et al. (2022), which generate answers by reformulating information across the context.
@@
Significant progress has been made in controllable text generation. Gururangan et al. (2020) fine-tune language models on domain-adaptive text to customize generated content attributes. Other methods include reinforcement learning Li et al. (2024), contrastive learning Zheng et al. (2023), and control codes for fine-tuning PLMs Keskar et al. (2019). Some approaches modify the probability distribution of PLMs, such as Liu et al. (2021) using two smaller “expert” models, and Yang and Klein (2021) conditioning generation with a “future discriminator.” Huang et al. (2023) explore multi-aspect text generation with trainable gates for enhanced control. Our proposed module, QASE, represents a novel adaptation of controlled text generation tailored to the specific challenges of MRC, with a focus on the precision and relevance of generated answers. Unlike methods that modify the overall generative process through complex architectural alterations or additional learning mechanisms, QASE directly utilizes the question and context to guide inferences.
@@ 4.1 Datasets and Metrics
We utilize three extractive MRC benchmark datasets:
@@ 4.2 Experimental Setup
4.3 Does QASE Mitigate Ill-Formed Generation?
@@
To assess QASE in mitigating ill-formed generation issue, we compare the performance of various PLMs fine-tuned with and without QASE, as detailed in Table 2. The conventional EM and partial match F1 scores effectively measure whether the generated answers match the gold answers in format on a token basis. Overall, models fine-tuned with QASE consistently outperform those without it when measured by overlap F1 score. Specifically, for the SQuAD dataset, models with QASE show an EM percentage increase of up to 33.8% and an F1 score improvement of up to 8.4% compared to vanilla fine-tuned models. For MultiSpanQA, improvements include up to 1.6% in EM F1 and up to 3.3% in overlap F1. Likewise, on the Quoref dataset, enhancements of up to 19.2% in EM percentage and up to 16.0% in F1 score are observed. These results confirm that QASE enables generative-based PLMs to produce more accurate, contextually coherent, and higher-quality answers in MRC tasks compared to vanilla fine-tuning approaches. We also include discussions on performance discrepancies across different datasets and base PLMs in Appendix B.3.
@@
4.3.1 QASE-Enhaced PLMs vs SOTA LLMs and Extractive Approaches
Our top model, Flan-T5-LargeQASE, is further benchmarked against leading models on each dataset’s official leaderboard, alongside zero-shot and few-shot GPT-3.5-Turbo and GPT-4. GPT-3.5-Turbo stands as one of OpenAI’s most efficient models in terms of capability and cost, while GPT-4 shows superior reasoning abilities Liu et al. (2023c). Studies indicate their superiority over traditional fine-tuning methods in most logical reasoning benchmarks Liu et al. (2023a). The prompts used to query the GPT variants in zero-shot and few-shot scenarios are detailed in Appendix A.4.
@@ 4.3.1 QASE-Enhaced PLMs vs SOTA LLMs and Extractive Approaches
On SQuAD, as showed in Table 3, Flan-T5-LargeQASE surpasses human performance, equaling the NLNet model from Microsoft Research Asia and the pre-trained BERT-Large Devlin et al. (2019). Additionally, it surpasses two-shot GPT-4 by 13.6% on EM and 4.0% on F1.
@@
On MultiSpanQA, Table 4 shows that Flan-T5-LargeQASE outperforms LIQUID Lee et al. (2023), which currently ranks #1 on the leaderboard, with respect to the overlap F1 score. Moreover, it surpasses zero-shot GPT-4 by 4.5% on the exact match F1 and 1.5% on the overlap F1, and two-shot GPT-4 by 2.3% on the exact match F1 and 0.8% on the overlap F1.
@@
On Quoref, Table 5 shows that Flan-T5-LargeQASE is comparable to CorefRoberta-Large Ye et al. (2020), which ranks #9 on the leaderboard, with a 0.5% higher exact match. Furthermore, it outperforms zero-shot GPT-4 by 11.9% on EM and 4.8% on F1, and two-shot GPT-4 by 2.5% on both EM and F1.
All top-performing models on these datasets’ leaderboards, equaling or exceeding Flan-T5-LargeQASE, are encoder-only extractive models. Therefore, these results demonstrate that QASE shortens or closes the gap between generative and extractive approaches, enhancing PLMs to match the capabilities of SOTA extractive models and outperform leading LLMs on extractive MRC.
4.4 Does QASE Improve Factual Consistency?
Flan-T5-Large $\bm{Q^{2}}$ F1 $\bm{Q^{2}}$ NLI
SQuAD no QASE 42.927 44.983
QASE 43.624 45.419
MultiSpanQA no QASE 32.889 31.433
QASE 34.732 36.452
4.5 Computational Cost
@@ 4.6 Ablation Studies
4.7 Qualitative Case Studies
In addition to the $Q^{2}$ statistical analysis in Section 4.4, we also perform qualitative case studies to further demonstrate the effectiveness of QASE in generating factual consistent answers.
@@ 5 Discussions
In this section, we briefly address the weak performance of Flan-T5 zero-shot and Llama 2 on extractive MRC tasks, despite their strong language understanding abilities. We note that a comprehensive analysis is beyond our study’s scope. Our goal is to gain insights into further improving these PLMs’ effectiveness in extractive MRC.
@@ 5.2 Llama 2 Performance
We also observe that models based on Llama 2 and Alpaca consistently underperform compared to those based on Flan-T5, across zero-shot and fine-tuned scenarios, with or without QASE. This discrepancy may arise from the significant difference in the number of trainable parameters, as illustrated in Table 1, during fine-tuning. Additionally, factors such as differences in pre-training datasets and varied adaptation to tasks due to structural disparities can also contribute to this performance gap. While acknowledging these factors, conducting a comprehensive comparison of different generative model architectures in extractive MRC tasks exceeds the scope of our study. For further discussion, please refer to Appendix B.2.
@@ 6 Conclusion and Future Work
In this study, we address out-of-control generation issue of generative PLMs in extractive MRC using QASE, a lightweight question-attended span extraction module, during the fine-tuning of PLMs. Our experiments show that QASE-enhanced PLMs generate better-quality responses with improved formality and factual consistency, matching SOTA extractive models and outperforming few-shot GPT-4 by a significant margin on all three extractive MRC datasets, bridging the gap between generative and extractive models in extractive MRC tasks. Importantly, QASE improves performance without a significant increase in computational costs, benefiting researchers with limited resources.
As the next step, we plan to conduct interpretability analyses to examine the performance discrepancies across different base PLMs and datasets.
In the future, we aim to evaluate our model on generative MRC tasks, such as Nguyen et al. (2016), to gauge its effectiveness in handling more intricate scenarios. Additionally, a significant emphasis will be placed on assessing the model’s overall capability in answer generation, with a specific focus on human perception. This involves incorporating human annotators alongside automatic metrics. Looking further ahead, we aspire to extend our research to explore strategies for mitigating input- and context-conflicting hallucinations in LLMs.
@@ A.4 Instruction Templates and Model Prompts
Table 13 provides the instruction and prompt templates used for fine-tuning the PLMs and for zero-shot and few-shot querying of PLMs and GPT variants across both single- and multi-span answer datasets. In few-shot prompting scenarios, examples are randomly selected from the training set.
@@
Zero-shot prompting PLMs and GPT variants on single-span answer dataset, SQuAD Instruction: Using the provided context, answer the question with exact phrases and avoid explanations. [Format the response as follows: ["answer1", "answer2", …].]
@@
Few-shot prompting PLMs and GPT variants Instruction: Using the provided context, answer the question with exact phrases and avoid explanations. [Format the response as follows: ["answer1", "answer2", …].]
Example i
Context: {example context}
Question: {example question}
Answer: example answer
@@
Table 13: Templates for fine-tuning instructions and zero-shot and few-shot query prompts. Text in square bracket is only added for multi-span answer datasets, MultiSpanQA and Quoref.
@@ B.2 Discussion on Llama 2 Performance
B.3 Discussion on Performance Discrepancy across Different Base PLMs and Datasets
Llama2 Alpaca Flan-T5 Small Flan-T5 Base Flan-T5 Large
$\Delta$EM
SQuAD 1.47 33.82 0.43 0.13 1.17
MultiSpanQA 1.61 -1.01 -0.08 0.32 -0.73
Quoref 19.24 - 4.28 3.30 1.36
$\Delta$F1
SQuAD 1.34 8.35 0.46 0.76 1.09
MultiSpanQA 3.30 1.32 0.80 0.11 1.36
Quoref 16.03 - 5.66 0.35 2.04
Table 15: Performance improvement (in %) of fine-tuned PLMs with QASE on each dataset.
As shown in Table 15, we observe a significant performance improvement with QASE across different base PLMs and datasets. Specifically, dataset-wise, a larger improvement is noted on Quoref compared to other datasets. This is partially due to the relatively weaker baseline performance on Quoref. For example, a fine-tuned Flan-T5-Large model without QASE achieves an F1 score of 90.71% on SQuAD, 83.09% on MultiSpanQA, and 80.49% on Quoref. Higher baseline scores indicate a strong initial performance, making further improvements more challenging and thus more meaningful. Despite the already high performance on the other two datasets, particularly SQuAD, the incorporation of QASE still results in noticeable improvements.
PLM-wise, we generally observe that the improvements on Llama2 and Alpaca are more substantial than those on the Flan-T5 base models, with few exceptions on MultiSpanQA. This trend can be partially attributed to the higher baseline performance of Flan-T5 models on these datasets. We discuss in Sections 5, B.1, and B.2 that factors such as (1) differences in pre-training datasets, with Flan-T5 models being fine-tuned on MRC tasks like SQuAD, and (2) varied adaptation to tasks due to structural disparities, can contribute to this performance gap. Encoder-decoder models, such as Flan-T5, are better equipped for tasks requiring extensive input processing, like MRC, making them more suitable for these tasks than decoder-only models, which are typically more suited to open-ended QA scenarios. This fundamental distinction partially accounts for Flan-T5’s superior performance in context-based question answering across both zero-shot and fine-tuned settings. While acknowledging these factors, a comprehensive comparison of different generative model architectures in MRC tasks exceeds the scope of our study.